# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v1

@@ Appendix B Experiment Details
For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting. We pre-train only with Shafto et al. [48], Taylor et al. [53]. We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects. We fine-tune with all data in Armeni et al. [2] and Gwilliams et al. [24] with the same train/val/test ratios as before.
@@ Appendix C Hyperparameters
Appendix D Compute Resources
All experiments were run on individual NVIDIA V100 GPUs with up to 32GiB of GPU memory on a system with up to 128GiB of RAM. Each pre-training run with the maximum amount of pre-training data took approximately 150 hours (6.25 days). Fine-tuning following pre-training took up to another 12 hours. We estimate that we used approximately 2000 hours of compute for the final experimental runs, including hyperparameter searches. In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
Appendix E Licences For Datasets And Code
@@
Appendix F Broader Impacts

### v2

@@ Appendix E Experiment Details
For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting. We pre-train only with Cam-CAN [48, 53]. We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects. We fine-tune with all data in Armeni et al. [2] and Gwilliams et al. [24] with the same train/val/test ratios as before.
@@ Appendix F Hyperparameters
Appendix G Compute Resources
All experiments were run on individual NVIDIA V100 and A100 GPUs with up to 40GiB of GPU memory on a system with up to 1TiB of RAM. Each pre-training run with the maximum amount of pre-training data took approximately 200 hours (8.3 days). Fine-tuning following pre-training took up to another 12 hours. We estimate that we used approximately 3000 hours of compute for the final experimental runs, including hyperparameter searches. In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
Appendix H Licences For Datasets And Code
@@
Appendix I Broader Impacts